# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

### v1

@@ D.3 Details of Performance Metrics
For real-world applications, the true Pareto fronts are typically unknown. For the convenience of comparing the performance of different algorithms, we utilize the approximate Pareto fronts provided by (Tanabe and Ishibuchi 2020) as the ground truth, which are generated with a large number of evaluations and have also been widely used in other related studies. To ensure fairness, we firstly generate the ideal point $\boldsymbol{y}_{\text{ideal}}$ and nadir point $\boldsymbol{y}_{\text{nadir}}$ from the approximate Pareto front of the tested problem:
@@
where $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}$ are solutions in approximate Pareto set $\mathcal{P}_{s}$ from the ground truth. After that, we normalize the resulting Pareto fronts obtained from the algorithms using the bounds of ideal point and nadir point. Finally, we set the reference point as $\boldsymbol{r}_{i}=1.1,i=\{1,2,...,m\}$ to calculate the hypervolume indicator of the normalized Pareto front obtained by each algorithm.
Appendix E Additional Experimental Results
E.1 Training Loss
@@
Figure 6: The mean loss convergences with the same data volume but different hyperparameter combinations.
We monitore the convergence of the mean loss during training presented by the model we used in this paper, represented by the blue solid line in Figure 6. Despite minor fluctuations, the loss curve shows that it has essentially converged. We further test the loss convergence by reducing the batch size and increasing the maximum sample length while maintaining the same amount of data. This adjustment results in the loss converging to a smaller level.
@@
E.2 Time Usage During Model Training
Figure 7: The proportion of time usage during model training (recorded every 10 epochs).
@@ E.2 Time Usage During Model Training
In our method, the data synthesis and model training are performed simultaneously. To evaluate the advantages of using synthetic data for pre-training, we monitored the proportion of time utilized during training, as illustrated in Figure 7. In the figure, ”data time” refers to the time spent synthesizing data for each epoch, ”forward time” indicates the duration spent on model inference to generate logits for loss calculation, and ”step time” denotes the time taken to compute the loss and perform gradient backpropagation to optimize the model parameters. Statistics results indicate that the model optimization and data synthesis account for the majority of pre-training time. Although a large amount of data is synthesized during training using our sampling mechanism, the time proportion spent on data synthesis is notably smaller than the time dedicated to optimizing model parameters. This efficiency can be largely attributed to the effective computation of the decomposition-based aggregation targets from extensive synthetic functions, which does not impose an unacceptable burden on pre-training as would be the case with expensive experimental data from real-world problems. Furthermore, it is noteworthy that a single forward propagation through the neural network is quite fast, accounting for only about 5% of the total time. This efficiency helps explain the rapid in-context inference observed once the pre-training is completed.
E.3 Dimensional Scalability
We further evaluate the scalability of our method in higher-dimensional spaces, including both feature and objective dimensions. We consider ZDT1-3 problems with a feature dimension of 30, as they readily allow for dimensional expansion in their problem formulations. In addition, we examine more challenging real-world problems with higher objective dimensions, including RE41 (car side impact design) and RE42 (conceptual marine design), both with 4 objectives, as well as RE61 (water resource planning), which has 6 objectives (Tanabe and Ishibuchi 2020). The experimental results are shown in Figure 8 and Table 4. Our method demonstrates strong adaptability and competitive performance on problems characterized by high-dimensional complex feature and objective spaces. Notably, the EI and UCB methods show clear advantages in synthetic problems. Despite being constrained by a small budget in the RE42 problem, their reported hypervolume metrics continue to improve compared to other methods. Although our UHVI performs slightly worse than EI and UCB, it remains competitive compared with the qParEGO method in most cases.
Figure 8: Sequential optimization performance of optimization metrics w.r.t. the number of observations (beyond initial points) on ZDT1-3 problems with 30 features, real-world applications RE41-42 with 4 objectives and RE61 with 6 objectives. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.
Table 4: Means (stds) of best optimization metrics in sequential optimization on ZDT1-3 problems with 30 features, real-world applications RE41-42 with 4 objectives and RE61 with 6 objectives over 10 independent runs. The best and second best mean results for each row are highlighted with shaded background and underlines respectively.
Test Instances Sobol qParEGO qEHVI qNEHVI FoMEMO-EI FoMEMO-UCB FoMEMO-UHVI
ZDT1 (30d2m) 2.271e+00(4.83e-02) 2.067e+00(1.19e-01) 1.341e+00(4.81e-02) 1.617e+00(7.89e-02) 1.202e+00(9.12e-02) 1.215e+00(9.11e-02) 1.896e+00(1.34e-01)
ZDT2 (30d2m) 3.541e+00(1.18e-01) 3.189e+00(9.29e-02) 2.418e+00(1.63e-01) 2.597e+00(1.51e-01) 2.175e+00(1.68e-01) 2.255e+00(1.72e-01) 3.512e+00(1.08e-01)
ZDT3 (30d2m) 2.021e+00(9.22e-02) 1.989e+00(8.86e-02) 1.866e+00(1.31e-01) 1.916e+00(9.50e-02) 1.629e+00(1.97e-01) 1.804e+00(1.18e-01) 2.082e+00(9.66e-02)
RE41 (7d4m) 5.462e-01(1.02e-02) 7.135e-01(1.41e-02) 7.905e-01(4.99e-03) 7.966e-01(5.92e-03) 7.503e-01(8.21e-03) 7.701e-01(9.51e-03) 7.148e-01(4.63e-03)
RE42 (6d4m) 2.328e-01(3.07e-02) 5.118e-01(4.33e-02) 7.850e-01(6.08e-03) 7.874e-01(6.08e-03) 7.554e-01(1.55e-02) 7.797e-01(1.37e-02) 6.224e-01(1.48e-02)
RE61 (3d6m) 1.192e+00(4.63e-02) 1.415e+00(2.20e-02) 1.441e+00(3.14e-02) 1.463e+00(6.98e-03) 1.461e+00(6.29e-03) 1.457e+00(1.52e-02) 1.387e+00(1.85e-02)
E.4 Parallel Scalability
We provide supplementary experimental results to further validate the scalability of our approach for parallel optimization. Results include the performance of all algorithms on synthetic problems with candidate batch sizes q=5 (shown in Figure 9 and Table 5) and q=10 (shown in Figure 10 and Table 6), as well as the performance of all algorithms on real-world applications with q=10 (shown in Figure 11 and Table 7). In addressing synthetic problems, our method demonstrates competitiveness with existing approaches, notably the UCB method for its robustness. Moreover, our method showcases scalability to high batch sizes of candidate solutions in various real-world applications, underscoring its potential value in diverse scenarios that allow for efficient batch experimentation. It is worth noting that the performance of our method diminishes as the batch size of candidates $q$ increases. This decline is attributed to our use of a relatively trivial parallelization strategy for searching batches of candidate solutions. More efficient batch search methods merit further exploration to enhance performance in larger batch scenarios.
Figure 9: Parallel optimization performance (q=5) of IGD+ metrics ($\downarrow$) w.r.t. the number of observations (beyond initial points) on synthetic problems. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.
Table 5: Means (stds) of best IGD+ metrics ($\downarrow$) in parallel optimization (q=5) on synthetic problems over 10 independent runs. The best and second best mean results for each row are highlighted with shaded background and underlines respectively.
Test Instances Sobol qParEGO qEHVI qNEHVI FoMEMO-EI FoMEMO-UCB FoMEMO-UHVI
ZDT1 1.466e+00(1.38e-01) 4.873e-01(1.55e-01) 1.686e-01(2.38e-02) 6.864e-02(1.48e-02) 7.055e-02(1.48e-02) 2.711e-02(5.30e-03) 5.046e-01(6.35e-02)
ZDT2 2.490e+00(1.33e-01) 8.501e-01(1.72e-01) 3.249e-01(9.56e-03) 3.350e-01(3.44e-17) 1.920e-01(6.87e-02) 3.497e-02(9.45e-03) 9.776e-01(1.02e-01)
ZDT3 1.386e+00(1.40e-01) 7.417e-01(1.25e-01) 4.544e-01(9.44e-02) 4.232e-01(9.79e-02) 4.152e-01(1.17e-01) 3.264e-01(7.84e-02) 9.348e-01(1.48e-01)
ZDT4 5.098e+01(6.19e+00) 3.953e+01(4.18e+00) 3.400e+01(4.24e+00) 3.523e+01(6.03e+00) 4.399e+01(3.84e+00) 4.079e+01(5.87e+00) 5.115e+01(3.08e+00)
ZDT6 6.638e+00(1.59e-01) 3.497e+00(1.38e+00) 1.615e-01(5.29e-02) 1.631e-01(6.78e-02) 5.363e-01(4.32e-01) 3.239e-01(2.81e-02) 2.919e+00(7.10e-01)
Omnitest 1.154e-01(1.39e-02) 6.341e-02(3.54e-02) 7.159e-02(3.31e-02) 1.170e-01(3.26e-02) 2.262e-02(1.78e-03) 1.970e-02(2.76e-03) 3.103e-02(3.65e-03)
Figure 10: Parallel optimization performance (q=10) of IGD+ metrics ($\downarrow$) w.r.t. the number of observations (beyond initial points) on synthetic problems. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.
Table 6: Means (stds) of best IGD+ metrics ($\downarrow$) in parallel optimization (q=10) on synthetic problems over 10 independent runs. The best mean result for each row is highlighted with shaded background.
Test Instances Sobol qParEGO qEHVI qNEHVI FoMEMO-EI FoMEMO-UCB FoMEMO-UHVI
ZDT1 1.360e+00(1.30e-01) 4.789e-01(1.67e-01) 9.813e-02(2.38e-02) 6.958e-02(1.60e-02) 6.052e-02(1.06e-02) 4.346e-02(1.35e-02) 5.686e-01(1.01e-01)
ZDT2 2.187e+00(2.32e-01) 6.831e-01(1.80e-01) 2.592e-01(4.67e-02) 3.355e-01(3.14e-03) 2.468e-01(1.01e-01) 5.722e-02(1.22e-02) 9.372e-01(1.68e-01)
ZDT3 1.523e+00(1.51e-01) 6.811e-01(9.58e-02) 4.376e-01(6.99e-02) 3.387e-01(6.37e-02) 5.666e-01(1.57e-01) 5.108e-01(1.62e-01) 8.398e-01(1.49e-01)
ZDT4 6.065e+01(4.73e+00) 3.774e+01(5.73e+00) 3.208e+01(4.53e+00) 3.878e+01(4.05e+00) 4.474e+01(3.42e+00) 4.508e+01(6.25e+00) 5.215e+01(4.47e+00)
ZDT6 6.544e+00(2.23e-01) 3.330e+00(1.42e+00) 1.199e-01(5.66e-02) 1.978e-01(7.44e-02) 1.145e+00(8.73e-01) 3.109e-01(4.75e-02) 2.160e+00(9.74e-01)
Omnitest 1.706e-01(1.46e-02) 6.888e-02(3.79e-02) 6.188e-02(2.08e-02) 1.101e-01(3.16e-02) 2.227e-02(3.86e-03) 2.061e-02(2.36e-03) 4.201e-02(5.41e-03)
Figure 11: Parallel optimization performance (q=10) of hypervolume metrics ($\uparrow$) w.r.t. the number of observations (beyond initial points) on real-world applications. The solid line is the mean value averaged over 10 independent runs, and the shaded region is the standard deviation with 95% confidence around the mean value.

### v2

@@ B.3 Performance Metrics
To rigorously evaluate the quality of the Pareto fronts obtained by different algorithms, we employ the widely-used hypervolume (HV) indicator (Zitzler and Thiele, 2002) as the performance metric. For both synthetic black-box functions and real-world hyperparameter optimization tasks, we adopt the evaluation protocol from BOFormer (Hung et al., 2025), in which the true objective values are normalized using the pre-defined bounds. Consequently, we set the reference point to $\boldsymbol{r}=[1.1,...,1.1]^{T}\in\mathbb{R}^{m}$ for HV computation in the normalized objective space. For real-world engineering design problems, we utilize the approximate Pareto fronts provided by (Tanabe and Ishibuchi, 2020) as the ground truth, which are generated with a large number of evaluations and have also been widely used in other related studies. To ensure fairness, we firstly generate the ideal point $\boldsymbol{y}_{\text{ideal}}$ and nadir point $\boldsymbol{y}_{\text{nadir}}$ from the approximate Pareto front of the tested problem:
@@
where $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}$ are solutions in approximate Pareto set $\mathcal{P}_{s}$ from the ground truth. After that, we normalize the resulting Pareto fronts obtained from the algorithms using the bounds of ideal point and nadir point. Finally, we also set the reference point as $\boldsymbol{r}=[1.1,...,1.1]^{T}\in\mathbb{R}^{m}$ to calculate the hypervolume indicator of the normalized Pareto front obtained by each algorithm.
Appendix C Analysis of Model Training and Ablation Study
In this section, we provide a more detailed analysis of the model training process. To further demonstrate the superiority of our selected model configuration, we compare the optimization results across different configurations. Here, “Base” refers to the model setup used throughout this paper, which employs an RBF kernel, an objective-aware regression head, and Tchebycheff aggregation. For ablation analysis, we retrain the model by altering one key component at a time while keeping the rest of the configurations unchanged. Specifically, we experiment with Matern32 and Matern52 kernels, a uniform regression head that is agnostic to the objective dimension, and a smoothed version of the Tchebycheff aggregation with a smoothing parameter of 0.05 (Lin et al., 2024b).
C.1 Training Loss
@@
Figure 4: The mean loss convergences with different model configurations.
We monitor the convergences of the mean loss during training presented by model with different configurations, represented in Figure 4. Despite minor fluctuations, each loss curve shows that it has essentially converged. We find that using the RBF kernel to generate data helps achieve a lower loss level, demonstrating the effectiveness of the RBF kernel in exploring a wider range of diverse data. While whether or not a objective-aware regression head is used has little impact on the reduction of loss. Notably, using smoothed aggregation leads to faster convergence and lower loss, which is related to the numerical stability brought by smoothing operation (Lin et al., 2024b), and thus facilitates gradient-based optimization of model parameters. However, there is a gap between the model pre-training phase and the subsequent downstream optimization process using the pre-trained models, and the performance of loss reduction cannot directly determine the optimization performance of the algorithm, which we will discuss further later.
C.2 Time Usage During Model Training
@@
Table 7: Averaged time usage per step during model training (in seconds).
Model Configuration Data Time Step Time Forward Time
Base 0.0474 0.0323 0.0065
Matern32 Kernel 0.0491 0.0325 0.0066
Matern52 Kernel 0.0489 0.0326 0.0067
Objective-Agnostic Head 0.0518 0.0325 0.0100
Smoothed Aggregation 0.0588 0.0382 0.0067
@@
In our method, the data synthesis and model training are performed simultaneously. To evaluate the advantages of using synthetic data for pre-training and in-context optimization with a single neural network forward pass, we record the detailed information on averaged time for each step utilized during training, as illustrated in Table 7. Among them, “data time” refers to the time spent synthesizing data for each step, “forward time” indicates the duration spent on model inference to generate logits for loss calculation, and “step time” denotes the time taken to compute the loss and perform gradient backpropagation to optimize the model parameters. Statistics results indicate that for all model configurations, the model optimization and data synthesis account for the majority of pre-training time. Although a large amount of data is synthesized during training using our sampling mechanism, the time spent on data synthesis is comparable to the time dedicated to optimizing model parameters. This efficiency can be largely attributed to the effective computation of the decomposition-based aggregation targets from extensive synthetic functions, which does not impose an unacceptable burden on pre-training as would be the case with expensive experimental data from real-world problems. Furthermore, it is noteworthy that a single forward propagation through the neural network is quite fast, this efficiency helps explain the rapid in-context inference observed once the pre-training is completed. In addition, we find that our chosen model configuration had advantages in all types of time usage.
C.3 Ablation Study of Optimization Performance
Table 8: Means (stds) of hypervolume metrics ($\uparrow$) on representative problems over 10 independent runs. The better mean result for each algorithm compared to the base configuration is marked in dark green.
Model Configuration Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
Base FoMEMO-EI 8.066e-01(6.2e-04) 7.308e-01(7.9e-03) 8.667e-01(2.0e-03) 9.107e-01(8.1e-03) 1.173e+00(2.0e-03) 1.261e+00(2.1e-02)
FoMEMO-UCB 8.075e-01(5.8e-04) 7.369e-01(8.5e-03) 8.722e-01(1.2e-03) 9.270e-01(4.5e-03) 1.175e+00(1.9e-03) 1.262e+00(1.8e-02)
FoMEMO-UHVI 8.052e-01(4.9e-04) 7.430e-01(8.4e-03) 8.609e-01(1.3e-03) 8.573e-01(9.6e-03) 1.176e+00(2.4e-03) 1.271e+00(1.6e-02)
FoMEMO-UR2I 8.016e-01(1.7e-03) 7.126e-01(9.1e-03) 8.741e-01(5.0e-04) 8.961e-01(1.1e-02) 1.179e+00(2.2e-03) 1.274e+00(1.2e-02)
Matern32 Kernel FoMEMO-EI 8.071e-01(5.0e-04) 7.354e-01(8.9e-03) 8.676e-01(1.1e-03) 9.134e-01(5.0e-03) 1.175e+00(2.0e-03) 1.263e+00(3.9e-03)
FoMEMO-UCB 8.063e-01(1.4e-03) 7.501e-01(1.1e-02) 8.693e-01(2.0e-03) 9.145e-01(8.4e-03) 1.175e+00(2.4e-03) 1.235e+00(2.6e-02)
FoMEMO-UHVI 7.841e-01(3.6e-02) 7.340e-01(7.8e-03) 8.560e-01(2.1e-03) 8.456e-01(9.8e-03) 1.175e+00(2.9e-03) 1.261e+00(2.1e-02)
FoMEMO-UR2I 8.018e-01(1.0e-03) 7.192e-01(8.6e-03) 8.697e-01(1.5e-03) 8.890e-01(7.8e-03) 1.174e+00(9.1e-03) 1.244e+00(1.6e-02)
Matern52 Kernel FoMEMO-EI 8.074e-01(6.0e-04) 7.430e-01(9.1e-03) 8.687e-01(9.9e-04) 9.090e-01(7.0e-03) 1.175e+00(2.6e-03) 1.254e+00(9.5e-03)
FoMEMO-UCB 8.083e-01(5.3e-04) 7.495e-01(9.1e-03) 8.736e-01(8.1e-04) 9.156e-01(6.5e-03) 1.177e+00(2.0e-03) 1.253e+00(2.0e-02)
FoMEMO-UHVI 8.074e-01(4.3e-04) 7.396e-01(7.6e-03) 8.610e-01(1.4e-03) 8.383e-01(8.0e-03) 1.175e+00(1.6e-03) 1.267e+00(1.4e-02)
FoMEMO-UR2I 7.975e-01(1.2e-03) 7.178e-01(1.1e-02) 8.753e-01(1.2e-03) 8.658e-01(1.1e-02) 1.178e+00(1.3e-03) 1.255e+00(6.7e-03)
Objective-Agnostic Head FoMEMO-EI 8.073e-01(9.1e-04) 7.192e-01(8.0e-03) 8.666e-01(1.9e-03) 9.105e-01(7.7e-03) 1.173e+00(2.5e-03) 1.261e+00(1.2e-02)
FoMEMO-UCB 8.085e-01(6.1e-04) 7.305e-01(9.2e-03) 8.721e-01(1.6e-03) 9.221e-01(4.6e-03) 1.177e+00(2.3e-03) 1.263e+00(1.2e-02)
FoMEMO-UHVI 8.070e-01(3.6e-04) 7.349e-01(1.1e-02) 8.639e-01(7.0e-04) 8.506e-01(9.9e-03) 1.174e+00(2.1e-03) 1.267e+00(1.6e-02)
FoMEMO-UR2I 8.069e-01(1.0e-03) 7.158e-01(6.8e-03) 8.769e-01(5.6e-04) 8.838e-01(6.1e-03) 1.176e+00(2.1e-03) 1.262e+00(9.1e-03)
Smoothed Aggregation FoMEMO-EI 7.995e-01(2.4e-03) 7.407e-01(7.0e-03) 8.629e-01(1.8e-03) 9.064e-01(7.5e-03) 1.176e+00(2.0e-03) 1.271e+00(1.6e-02)
FoMEMO-UCB 8.027e-01(1.4e-03) 7.397e-01(8.9e-03) 8.683e-01(1.8e-03) 9.279e-01(3.5e-03) 1.177e+00(1.7e-03) 1.269e+00(1.3e-02)
FoMEMO-UHVI 7.977e-01(1.4e-03) 7.280e-01(8.8e-03) 8.577e-01(3.3e-03) 8.329e-01(1.7e-02) 1.174e+00(1.9e-03) 1.266e+00(1.6e-02)
FoMEMO-UR2I 7.903e-01(1.9e-03) 6.897e-01(5.0e-02) 8.669e-01(7.7e-04) 8.857e-01(4.6e-03) 1.176e+00(1.9e-03) 1.267e+00(1.3e-02)
We further compared the optimization performance of all proposed acquisition functions with different model configurations on 6 representative problems, as shown in Table 8. For each algorithm with different model configurations, if it outperforms the corresponding algorithm in base configuration, we mark it in dark green to visualize the impact of different model configurations on the final optimization performance. The results show that our chosen model configuration performs better in most cases. We find that using Matern32 and Matern52 yield better results especially for synthetic problems involving Branin functions. The reason may lie that the RBF kernel, being infinitely differentiable, imposes a strong smoothness prior that hinders its ability to capture the localized transitions and high-curvature regions between the peaks and valleys of the Branin landscape. In contrast, the Matern class of kernels relaxes these smoothness constraints, which effectively balances the global trend capturing and local flexibility required to navigate the distinctive landscape of the Branin function. In addition, we observe that using the objective-agnostic head yields inferior results compared to the base configuration in most cases. This demonstrates that while loss gaps may be marginal during training, the adaptive selection of dimension-specific regressors during inference is pivotal for enhancing final optimization outcomes. Finally, empirical results indicate that the configuration with smoothed aggregation generally underperforms the baseline. This disparity likely stems from the fact that while smoothing facilitates model training, the approximation errors introduced by smoothing operation may bias the accurate fitting of the optimal region in the exact aggregation function (Lin et al., 2024b), further misleading the precise identification of the global optimum in the sub-problem during the optimization process.
Appendix D Analysis of Method Scalability
D.1 Dimensional Scalability